OpenStudy (anonymous):

quite urgent, help appreciated n pellet Let f(x) = (3x^3)/((4-5x)^5) find the equation of the line tangent to the graph of f at x=2

OpenStudy (anonymous):

differentiate f(x) \[f'(x)=(4-5x)^5\times9x^2-3x^3\times5(4-5x)^4\times-5, all divided by the \square of (4-5x)^5\]

OpenStudy (anonymous):

sub x=2 into expression for f'(x) to get the slope m of the tangent. Then sub x=2 into f(x) to find the y-coordinate Then line tangent equation is \[y-y_1=m(x-x_1)\]